OpenStudy (abbie):

10 medals!! please help?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you would need 10 accounts to give 10 medals

OpenStudy (noseboy908):

Nonetheless, I'd be willing to help if I'm able.

OpenStudy (ankit042):

What is the question?

OpenStudy (abbie):

this is what i do -- I go to your profile and give you medals for previous questions you have answered....

OpenStudy (anonymous):

ill give you if you really want? :))

OpenStudy (abbie):

@ankit @noseboy @niko can someone help?

OpenStudy (abbie):

@noseboy908

hartnn (hartnn):

Hi abbie, you can use the fact that Volume of a solid is proportional to cube of the side! and Surface Area of the solid is proportional to square of the side!

hartnn (hartnn):

So, Since we are given volume of both the solids, lets first find the ratio of sides, \(\Large \dfrac{V_1}{V_2}= \dfrac{s_1^3}{s_2^3}\) \(\Large \dfrac{1331}{216}= \dfrac{s_1}{s_2}\) take cube root on both sides, and you will get the ratio of sides!

hartnn (hartnn):

that is supposed to be \(\Large \dfrac{1331}{216}= \dfrac{s_1^3}{s_2^3}\)

hartnn (hartnn):

know whats the cube root of 1331 and 216 ?

OpenStudy (abbie):

I dont get it... so what do I have to do?

hartnn (hartnn):

take the cube root on both sides of \(\Large \dfrac{1331}{216}= \dfrac{s_1^3}{s_2^3}\)

hartnn (hartnn):

\(\Large \dfrac{\sqrt[3]{1331}}{\sqrt[3]{216}}= \dfrac{s_1}{s_2}\) so you will have to find out the cube root of 1331 and 216

OpenStudy (abbie):

11/6

hartnn (hartnn):

correct! :)

hartnn (hartnn):

so, the ratio of sides is s1/s2 = 11/6 now since the surface area is proportional to square of sides, we have \(\Large \dfrac{SA_1}{SA_2} = (\dfrac{s_1}{s_2})^2 \) \(\Large \dfrac{484}{SA_2} = (\dfrac{11}{6})^2 \) just find SA2 from this !

OpenStudy (abbie):

what next?

hartnn (hartnn):

use 11^2 = 121 6^2 = 36

hartnn (hartnn):

then cross-multiply!

OpenStudy (abbie):

I got 144

hartnn (hartnn):

and you're absolutely correct! :)

OpenStudy (abbie):

Thankyou so much ^_^

hartnn (hartnn):

welcome so much ^_^